OpenStudy (anonymous):

to determine if ABCD is a parallelogram what is the midpoint of BD and AC? A(-5,-5) B(-2,4) C (6,4) D (3,-5)

OpenStudy (angela210793):

M is the midpoint of AC Xm=(Xa+Xa)/2=(-5+6)/2=1/2 and Ym=(Ya+Yc)/2=(-5+4)/2=-1/2

OpenStudy (angela210793):

Same for BD

OpenStudy (angela210793):

Get it?

OpenStudy (anonymous):

so its 1/2?

OpenStudy (angela210793):

It's the point M(1/2;-1/2)

OpenStudy (angela210793):

And you do the same For BD just that u replace the coordinates of B and D